OpenStudy (anonymous):

given that (9,2) is on the graph of f (x) find the corresponding point for the function f (-3/4 x)

OpenStudy (tkhunny):

We know that f(9) = 2. Does that tell us anything about f(-3/4 x)?

OpenStudy (anonymous):

i wrote down everything the question asked

OpenStudy (tkhunny):

Well, it makes no sense. What class are you in? What section of the book (if you have one)? OR maybe the name fo the unit you are working on?

OpenStudy (anonymous):

functional values, its algebra 2 on a computer class.

OpenStudy (tkhunny):

Well, if -¾x = 9, we have x = -12 and f((-3/4)(-12)) = f(9) = 2 I see. We have a 25% expansion and a reflection across the y-axis. Ringing any bells?

OpenStudy (anonymous):

the answer is 2? i really just dont understand this lesson

OpenStudy (tkhunny):

Like I said, I really don't know what it's asking. If we are to compare f(x) to f(-¾x), I'm not sure what "2" helps us see.

OpenStudy (anonymous):

Given that (-9,3) is on the graph pf f (x) find the corresponding point for the function f (-3x)

OpenStudy (anonymous):

can you help me please???

OpenStudy (tkhunny):

That's a compression factor fo 3 and a reflection across the y-axis. Do you have even one solved example? I'm just guessing what it wants. Perhaps the first one was (-12,2) and this one is (3,3).

OpenStudy (anonymous):

given that (-2,-1) is on the graph of f (x) find the corresponding point for the function f (x-4)

OpenStudy (tkhunny):

Find the value of x that gives the new function's argument as -2. x-4 = -2 This is a translation.

OpenStudy (anonymous):

i think

OpenStudy (tkhunny):

What new value for x gives the new argument a functional value of -2

OpenStudy (anonymous):

given that (5,-4) is on the graph of f(x) find the corresponding point for the function -1/4 f(x)

OpenStudy (anonymous):

hello

OpenStudy (tkhunny):

I'm tired of doing these. It's your turn to show your work. Get the last one right before you move on to this one.